# Adversarial generalization of unfolding

### v1

@@ 1 Introduction
Table 1: Comparison of our adversarial generalization error bounds to related work’s. Bold letters indicate the method yielding the tighter – which is desirable – upper-bounds, in terms of attack level $\varepsilon$ and number of layers $L$. For [12, 49], $s_{k}$ denotes the upper bound on the spectral norm of the parameter matrix at the $k$th layer output.
@@
Adv. gen. error bounds $\mathcal{O}(\varepsilon\prod_{k=1}^{L}s_{k})$ $\mathcal{O}(\varepsilon\prod_{k=1}^{L}s_{k})$ $\mathcal{O}(\sqrt{L\log(\varepsilon)})$
@@ 2 Background and problem formulation
Notation. For matrices $A_{1},A_{2}\in\mathbb{R}^{N\times N}$, we denote by $A_{1};A_{2}]\in\mathbb{R}^{2N\times N}$ their concatenation with respect to the first dimension, and $A_{1}\,|\,A_{2}]\in\mathbb{R}^{N\times 2N}$ their concatenation with respect to the second dimension. We write $O_{n\times n}\in\mathbb{R}^{n}$ for the zero matrix and $I_{n\times n}$ for the $n\times n$ identity matrix. For $x\in\mathbb{R},\,\tau>0$, the soft thresholding operator $\mathcal{S}_{\tau}:\mathbb{R}\mapsto\mathbb{R}$ is defined as $\mathcal{S}_{\tau}(x)=\mathrm{sign}(x)\max(0,|x|-\tau)$. For $x\in\mathbb{R}^{n}$, $\mathcal{S}_{\tau}(\cdot)$ acts component-wise and is 1-Lipschitz with respect to $x$. The covering number $\mathcal{N}(T,d,t)$ of a metric space $(T,d)$ at level $t>0$, is defined as the smallest number of balls with respect to the metric $d$ required to cover $T$. When $d$ is induced by some norm $\|\cdot\|$, we write $\mathcal{N}(T,\|\cdot\|,t)$.
@@
where $T_{W}(u)=\Lambda_{W}u+(A^{T}A+\rho W^{T}W)^{-1}A^{T}y\in\mathbb{R}^{n}$, with $\Lambda_{W}=[-\rho(A^{T}A+\rho W^{T}W)^{-1}W^{T}\,|\,\rho(A^{T}A+\rho W^{T}W)^{-1}W^{T}]\in\mathbb{R}^{n\times 2N}$. More details on ADMM and ADMM-DAD can be found at Appendix LABEL:unfoldparticulars.
@@ 3 Main results
We address this paper’s research question ($\star$1), by delivering adversarial generalization error bounds for ADMM-DAD in the form of Theorem 3 (with proof found in Appendix LABEL:genapp) and Corollary 4.
@@
Let $\tilde{h}^{L}_{W}(\cdot)$ be the perturbed final decoder implemented by ADMM-DAD and defined in (7), $L\geq 2$, and learnable overcomplete sparsifier $W\in\mathcal{F}_{\beta}$. Then, for any $W_{1},\,W_{2}\in\mathcal{F}_{\beta}$, we have $\begin{split}\|\tilde{h}^{L}_{W_{1}}(Y)-\tilde{h}^{L}_{W_{2}}(Y)\|_{F}:=\|h^{L}_{W_{1}}(Y+\Delta_{1})-h^{L}_{W_{2}}(Y+\Delta_{2})\|_{F}\leq\mathrm{Lip}_{h}^{L,\varepsilon}\|W_{1}-W_{2}\|_{2\rightarrow 2},\end{split}$ (13) where $\Delta_{i}:=\Delta_{W_{i}}^{\text{FGSM}}$ with $\|\Delta_{W_{i}}^{\text{FGSM}}\|_{F}\leq\sqrt{s}\varepsilon$, $i=1,2$, and Lipschitz constants $\mathrm{Lip}_{h}^{L,\varepsilon}$ calculated explicitly (see (LABEL:applip) of Appendix LABEL:applipdec), and depending exponentially on the number of layers $L$ and linearly on the attack level $\varepsilon$.
@@
Why is Theorem 8 important? The ARC is an essential tool for thoroughly explaining adversarial generalization. The explicit dependence of the ARC estimate on elemental quantities like $N$, $L$, $\varepsilon$, stresses how these ripple out to the adversarial generalization of ADMM-DAD. Especially, by definition of $\mathrm{Lip}_{h}^{L,\varepsilon}$, due to the appearance of the logarithm, and since (16) can be proven to be integral-free (cf. Appendix LABEL:genapp), the ARC estimate roughly scales like $\sqrt{NL\log(\varepsilon)}$ (cf. Corollary 4).
@@
We combine (17) and [38, Theorem 26.5], to give adversarial generalization error bounds for ADMM-DAD, stated formally in Theorem 3 and Corollary 4, thus answering this paper’s research question ($\star$1). The proofs of Proposition 5, Theorem 6, Proposition 7 and Theorem 8 can be found at Appendices LABEL:appbounded, LABEL:applipdec, LABEL:coverapp and LABEL:arcapp, respectively. Moreover, [11, Theorem 8.23], [38, Theorem 26.5] and [28, Corollary 4] are formally stated as Theorem LABEL:dudleyapp, Theorem LABEL:radem and Lemma LABEL:contraction, respectively. In the next Section, we assess the validity of our theory with a series of experiments on real-world data.
Figure 1: Performance of ADMM-DAD on CIFAR10 (left) and SVHN (right), for varying attack levels $\varepsilon$ and number of layers $L$, and overcompleteness $N=10n$. Top: clean test MSE (18) and adversarial test MSE (19). Bottom: adversarial EGE (20). For both datasets, (20) increases as Theorem 3 suggests, and in fact scales at the rate dictated by Corollary 4, thus confirming our derived generalization theory. A similar increment is observed for both (18) and (19), but at a reasonable rate, thereby highlighting the adversarial robustness of ADMM-DAD.
@@ 4 Experiments
We train and test ADMM-DAD on two real-world image datasets: CIFAR10 (50000 training and 10000 test $32\times 32$ coloured image examples) and SVHN (73257 training and 26032 test $32\times 32$ colored image examples). For both datasets, we transform the images into grayscale ones and vectorize them. We fix $m/n=25\%$, and alternate the overcompleteness $N$, and the number of layers $L$. We consider a standard CS setup, with an appropriately normalized random Gaussian $A\in\mathbb{R}^{m\times n}$, and noisy observations of the form $y=Ax+e$, with $e$ being zero-mean Gaussian noise, with standard deviation $\mathrm{std}=10^{-2}$. To generate the adversarial attack $\delta$, we employ FGSM from [15], under an $l_{2}$-norm constraint, and attack ADMM-DAD with $\delta$ during training and test times, with varying attack levels $\varepsilon$. We initialize the learnable overcomplete sparsifier $W\in\mathbb{R}^{N\times n}$ using a Xavier normal distribution [14]. We implement all models in PyTorch [19] and train them using the Adam algorithm [21], with batch size $128$. As evaluation metrics, we use the clean and adversarial test MSEs
@@
with $\widetilde{\mathcal{L}}_{\text{train}}(h)$ being the adversarial train MSE defined in Sec. 2. We also compare ADMM-DAD to a state-of-the-art baseline DUN called ISTA-net [4], parameterized by an orthogonal sparsifier. We aim to use the structural difference between the two DUNs to showcase how the adversarial robustness and generalization of DUNs are affected, when employing an overcomplete sparsifier instead of an orthogonal one. For more experimental settings and details, we refer the reader to Appendix LABEL:expdetails.
@@
Test and generalization errors for increasing $\varepsilon$. We measure the performance of ADMM-DAD with $L=5$ and $L=10$ on CIFAR10, and $L=10$ and $L=15$ on SVHN, both with fixed $N=10n$, in terms of the clean test MSE (18), the adversarial test MSE (19), and the adversarial EGE (20), as $\varepsilon$ varies. We report the results in Figure 1, corroborating Theorem 3. Specifically, for both datasets, we observe that the adversarial EGEs, as these are depicted in the bottom of Figure 1, increase, as both $L$ and $\varepsilon$ increase. Despite the appearance of other terms in (10), the adversarial EGEs seem to scale at the rate of $\sqrt{L\log(\varepsilon)}$, like Corollary 4 suggests. As illustrated at the top of Figure 1, both (18) and (19) also increase as $\varepsilon$ increases. This behavior seems reasonable from the adversarial robustness perspective, since the higher the attack level, the more a neural network “struggles” to infer correctly. Nevertheless, the increment of (18) and (19) on both datasets seem to be at the reasonable rate of roughly square-root. Overall, the empirical behavior of ADMM-DAD matches its theoretical one.
@@
Figure 2: Robustness plots of 5-layer ADMM-DAD on CIFAR10, (left) and 10-layer ADMM-DAD on SVHN (right), for alternating overcompleteness $N$ and different attack levels $\varepsilon$. For both datasets, as $N$ increases, the clean test MSEs (18) and the adversarial test MSEs (19) drop. Importantly, for the standard case of $\varepsilon=0.01$, the robustness gap of ADMM-DAD on both datasets is particularly small. All in all, results highlight the beneficial role that NN plays on robustifying ADMM-DAD against varying adversarial attack levels.
@@
The role of $N$ in adversarial robustness. We measure the adversarial robustness of 5- and 10-layer ADMM-DAD on CIFAR10 and SVHN, respectively, for increasing $N$, and three different values of $\varepsilon$, by means of (18) and (19). Given that DUNs operate in a regression setting (see (7)) – so that standard classification-type metrics like accuracy cannot be utilized – the deployment of the clean and adversarial test MSEs (and their gap) as adversarial robustness metrics is a common practice [35, 36]. We report the results for both datasets in Figure 2, which demonstrates an intriguing phenomenon: as $N$ increases, for a fixed $\varepsilon$, both clean and adversarial test MSEs drop, indicating that the overcompleteness of the learnable sparsifying transform promotes ADMM-DAD’s adversarial robustness. Of course, the MSEs increase as $\varepsilon$ also increases, but this is anticipated, since a stronger attack in the CS observations tantamounts to a DUN being less able to recover the data. Nevertheless, explaining the adversarial robustness of ADMM-DAD through the overcompleteness of the learnable sparsifier lays a fruitful ground, to identify “hyper-parameters” and properly fine-tune them, so as to boost adversarial robustness in the unfolding regime (see also Sec. 5).
@@
A note on the robustness gap. Based on Figures 1 and 2, we examine the robustness gap of ADMM-DAD, i.e., the difference between (18) and (19). We notice that the robustness gap slightly increases for the CIFAR10 with increasing $\varepsilon$ – albeit the gap’s scaling from one attack level to the next is reasonable – but the picture, e.g., Figure 1, is much better for the SVHN, where both test errors grow proportionally. This indicates that ADMM-DAD enjoys adversarial robustness to increasing attack levels. Especially in the case of $\varepsilon=0.01$ – which is still a non-negligible attack level – we deduce from Figure 2 that, for both datasets, the robustness gap is especially small, e.g., $\sim 10^{-4}$. Due to these interesting findings, explaining the robustness gap of DUNs, both from a theoretical and practical side, could be an inspiring future work (see also Sec. 5).
@@
Table 2: Comparison of ADMM-DAD – with overcompleteness $N=10n$ – to the baseline, both with 10 layers, against different attack levels $\varepsilon$, on CIFAR10 (top) and SVHN (bottom). Bold letters indicate the DUN that scores the best performance in terms of all metrics (18), (19), (20). Overall, ADMM-DAD outperforms the baseline, highlighting the advantage of overcompleteness over orthogonality in the unfolding regime.

### v3

@@ 1 Introduction
Overall, our proposed study provides the first theoretical framework for understanding and improving the adversarial robustness of DUNs in CS, with direct implications for safety-critical applications. In medical imaging, for example, our results could help ensure that MRI reconstructions remain reliable even under adversarial perturbations, reducing the risk of misdiagnosis. Furthermore, by showcasing that overparameterization via learnable overcomplete sparsifiers improves robustness, our work offers concrete design principles for developing resilient, interpretable networks in real-world deployments, where accuracy, efficiency, and security must coexist.
2 Related work
@@
Table 1: Comparison of our adversarial generalization error bounds to related work’s. Bold letters indicate the method yielding the tighter – which is desirable – upper-bounds, in terms of attack level $\varepsilon$ and number of layers $L$. For [12, 49], $\beta_{k}$ denotes the upper bound on the spectral norm of the parameter matrix at the $k$th layer output. For our work, the upper bound on the spectral norm of the parameter matrix shared across all layers is $\sqrt{\beta}$.
@@
Adv. gen. error bounds $\mathcal{O}(\varepsilon\prod_{k=1}^{L}\beta_{k})$ $\mathcal{O}(\varepsilon\prod_{k=1}^{L}\beta_{k})$ $\mathcal{O}(\sqrt{L\log(\beta\varepsilon)})$
@@ 3 Background and problem formulation
Notation. For matrices $A_{1},A_{2}\in\mathbb{R}^{N\times N}$, we denote by $A_{1};A_{2}]\in\mathbb{R}^{2N\times N}$ their concatenation with respect to the first dimension, and $A_{1}\,|\,A_{2}]\in\mathbb{R}^{N\times 2N}$ their concatenation with respect to the second dimension. We write $O_{n\times n}\in\mathbb{R}^{n\times n}$ for the zero matrix and $I_{n\times n}$ for the $n\times n$ identity matrix. For $x\in\mathbb{R},\,\tau>0$, the soft thresholding operator $\mathcal{S}_{\tau}:\mathbb{R}\mapsto\mathbb{R}$ is defined as $\mathcal{S}_{\tau}(x)=\mathrm{sign}(x)\max(0,|x|-\tau)$. For $x\in\mathbb{R}^{n}$, $\mathcal{S}_{\tau}(\cdot)$ acts component-wise and is 1-Lipschitz with respect to $x$. The covering number $\mathcal{N}(T,d,t)$ of a metric space $(T,d)$ at level $t>0$, is defined as the smallest number of balls with respect to the metric $d$ required to cover $T$. When $d$ is induced by some norm $\|\cdot\|$, we write $\mathcal{N}(T,\|\cdot\|,t)$.
@@
where $T_{W}(u)=\Lambda_{W}u+(A^{T}A+\rho W^{T}W)^{-1}A^{T}y\in\mathbb{R}^{n}$, with $\Lambda_{W}=[-\rho(A^{T}A+\rho W^{T}W)^{-1}W^{T}\,|\,\rho(A^{T}A+\rho W^{T}W)^{-1}W^{T}]\in\mathbb{R}^{n\times 2N}$. More details on ADMM and ADMM-DAD can be found at Appendix A.
@@ 4 Main results
We address this paper’s research question ($\star$1), by delivering adversarial generalization error bounds for ADMM-DAD in the form of Theorem 3 (with proof found in Appendix C.5) and Corollary 4.
@@
Let $\tilde{h}^{L}_{W}(\cdot)$ be the perturbed final decoder implemented by ADMM-DAD and defined in (7), $L\geq 2$, and learnable overcomplete sparsifier $W\in\mathcal{F}_{\beta}$. Then, for any $W_{1},\,W_{2}\in\mathcal{F}_{\beta}$, we have $\begin{split}\|\tilde{h}^{L}_{W_{1}}(Y)-\tilde{h}^{L}_{W_{2}}(Y)\|_{F}:=\|h^{L}_{W_{1}}(Y+\Delta_{1})-h^{L}_{W_{2}}(Y+\Delta_{2})\|_{F}\leq\mathrm{Lip}_{h}^{L,\varepsilon}\|W_{1}-W_{2}\|_{2\rightarrow 2},\end{split}$ (13) where $\Delta_{i}:=\Delta_{W_{i}}^{\text{FGSM}}$ with $\|\Delta_{W_{i}}^{\text{FGSM}}\|_{F}\leq\sqrt{s}\varepsilon$, $i=1,2$, and Lipschitz constants $\mathrm{Lip}_{h}^{L,\varepsilon}$ depending exponentially on the number of layers $L$ and linearly on the attack level $\varepsilon$: $\begin{split}\mathrm{Lip}_{h}^{L,\varepsilon}=&2\gamma\rho\sqrt{\beta}\Bigg((r\nu)^{L-1}\gamma\|A\|_{2\rightarrow 2}\bigg(r\|Y\|_{F}+r\sqrt{s}\varepsilon+2\beta(B_{\mathrm{in}}+B_{\mathrm{out}})^{2}\frac{\sqrt{s}\varepsilon}{\kappa^{2}}\nu\gamma^{2}\|A\|_{2\rightarrow 2}^{2}\bigg)\\ &+\sum_{k=2}^{L}(r\nu)^{L-k}H_{k}+\nu^{2}\gamma\|A\|_{2\rightarrow 2}(\|Y\|_{F}+\sqrt{s}\varepsilon)\bigg(1+\sum_{k=1}^{L-1}(r\nu)^{k}\bigg)\Bigg),\end{split}$ (14) with $r=1+2\beta\gamma\rho$, $\gamma$ as in Proposition 5, $\beta$ as in Definition 1, $\nu=(1+\sqrt{2})$, and $H_{k}$ a constant calculated explicitly and defined in (49) of Appendix C.2.
@@
Why is Theorem 8 important? The ARC is an essential tool for thoroughly explaining adversarial generalization. The explicit dependence of the ARC estimate on elemental quantities like $N$, $L$, $\varepsilon$, stresses how these ripple out to the adversarial generalization of ADMM-DAD. Especially, by definition of $\mathrm{Lip}_{h}^{L,\varepsilon}$, due to the appearance of the logarithm, and since (17) can be proven to be integral-free (cf. Appendix C.5), the ARC estimate roughly scales like $\sqrt{NL\log(\varepsilon)}$ (cf. Corollary 4).
@@
We combine (18) and [38, Theorem 26.5], to give adversarial generalization error bounds for ADMM-DAD, stated formally in Theorem 3 and Corollary 4, thus answering this paper’s research question ($\star$1). The proofs of Proposition 5, Theorem 6, Proposition 7 and Theorem 8 can be found at Appendices C.1, C.2, C.3 and C.4, respectively. Moreover, [11, Theorem 8.23], [38, Theorem 26.5] and [28, Corollary 4] are formally stated as Theorem B.6, Theorem B.7 and Lemma B.8, respectively. In the next Section, we assess the validity of our theory with a series of experiments on real-world data.
@@ 5 Experiments
We train and test ADMM-DAD on two real-world image datasets: CIFAR10 (50000 training and 10000 test $32\times 32$ coloured image examples) and SVHN (73257 training and 26032 test $32\times 32$ colored image examples). For both datasets, we transform the images into grayscale ones and vectorize them. We fix $m/n=25\%$, and alternate the overcompleteness $N$, and the number of layers $L$. We consider a standard CS setup, with an appropriately normalized random Gaussian $A\in\mathbb{R}^{m\times n}$, and noisy observations of the form $y=Ax+e$, with $e$ being zero-mean Gaussian noise, with standard deviation $\mathrm{std}=10^{-2}$. To generate the adversarial attack $\delta$, we employ FGSM from [15], under an $l_{2}$-norm constraint, and attack ADMM-DAD with $\delta$ during training and test times, with varying attack levels $\varepsilon$. To highlight the adversarial robustness of ADMM-DAD against more powerful attacks, we also employ [15] to generate an $\ell_{2}$-based PGD attack with 10 iterations. We initialize the learnable overcomplete sparsifier $W\in\mathbb{R}^{N\times n}$ using a Xavier normal distribution [14]. We implement all models in PyTorch [19] and train them using the Adam algorithm [21], with batch size $128$. As evaluation metrics, we use the clean and adversarial test MSEs
@@
with $\widetilde{\mathcal{L}}_{\text{train}}(h)$ being the adversarial train MSE defined in Sec. 3. We also compare ADMM-DAD to a state-of-the-art baseline DUN called ISTA-net [4], parameterized by an orthogonal sparsifier. We aim to use the structural difference between the two DUNs to showcase how the adversarial robustness and generalization of DUNs are affected, when employing an overcomplete sparsifier instead of an orthogonal one. For more experimental settings and details, we refer the reader to Appendix D.
Figure 1: Performance of ADMM-DAD (plotted on logarithmic scale) on CIFAR10 (left) and SVHN (right), for varying number of layers $L$ and attack levels $\varepsilon$ of the FGSM, and overcompleteness $N=10n$. Top: clean test MSE (19) and adversarial test MSE (20). Bottom: adversarial EGE (21). For both datasets, (21) increases as Theorem 3 suggests, and in fact scales at the rate dictated by Corollary 4, thus confirming our derived generalization theory. A similar increment is observed for both (19) and (20), but at a reasonable rate, thereby highlighting the adversarial robustness of ADMM-DAD.
Figure 2: Performance of ADMM-DAD (plotted on logarithmic scale) on CIFAR10 (left) and SVHN (right), for varying number of layers $L$ and attack levels $\varepsilon$ of the PGD (10 iterations), and overcompleteness $N=10n$. Top: clean test MSE (19) and adversarial test MSE (20). Bottom: adversarial EGE (21). Although our theoretical analysis focuses on FGSM, we observe that even for a stronger adversarial attack like PGD, (21) scales at the rate dictated by Corollary 4, for both datasets, thus corroborating our derived generalization theory. Similarly, (19) and (20) also increase, but at a reasonable rate, thus highlighting the adversarial robustness of ADMM-DAD, even under more powerful than FGSM attacks.
@@
Test and generalization errors for increasing $\varepsilon$. We measure the performance of ADMM-DAD with $L=5$ and $L=10$ on CIFAR10, and $L=10$ and $L=15$ on SVHN, both with fixed $N=10n$, in terms of the clean test MSE (19), the adversarial test MSE (20), and the adversarial EGE (21), as $\varepsilon$ varies, for both the FGSM and the PGD. We report the results in Figure 1 for FGSM, and in Figure 2 for PGD, both corroborating Theorem 3. Specifically, for both datasets and attacking methods, we observe that the adversarial EGEs, as these are depicted in the bottom of Figure 1 and Figure 2, increase, as both $L$ and $\varepsilon$ increase. Despite the appearance of other terms in (10), the adversarial EGEs seem to scale at the rate of $\sqrt{L\log(\varepsilon)}$, like Corollary 4 suggests. As illustrated at the top of Figure 1 and Figure 2, both (19) and (20) also increase as $\varepsilon$ increases. This behavior is anticipated from the adversarial robustness perspective, since the higher the attack level, the more a neural network “struggles” to infer correctly. Nevertheless, the increment of (19) and (20) on both datasets seem to be at the reasonable rate of roughly square-root. Overall, the empirical behavior of ADMM-DAD matches its theoretical one. Interestingly, despite the fact that our theory hinges upon the FGSM, the similar adversarial robustness and generalization of the PGD-attacked ADMM-DAD fuels us to theoretically investigate this phenomenon more in the future, as we pinpoint in Sec. 6.
@@
Figure 3: Robustness plots of 5-layer ADMM-DAD on CIFAR10, (left) and 10-layer ADMM-DAD on SVHN (right), for alternating overcompleteness $N$ and different attack levels $\varepsilon$ of FGSM. For both datasets, as $N$ increases, the clean test MSEs (19) and the adversarial test MSEs (20) drop. Importantly, for the standard case of $\varepsilon=0.01$, the robustness gap of ADMM-DAD on both datasets is particularly small. All in all, results highlight the beneficial role that NN plays on robustifying ADMM-DAD against varying adversarial attack levels.
@@
The role of $N$ in adversarial robustness. We measure the adversarial robustness of 5- and 10-layer ADMM-DAD on CIFAR10 and SVHN, respectively, for increasing $N$, and three different values of $\varepsilon$ for the FGSM, by means of (19) and (20). Given that DUNs operate in a regression setting (see (7)) – so that standard classification-type metrics like accuracy cannot be utilized – the deployment of the clean and adversarial test MSEs (and their gap) as adversarial robustness metrics is a common practice [35, 36]. We report the results for both datasets in Figure 3, which demonstrates an intriguing phenomenon: as $N$ increases, for a fixed $\varepsilon$, both clean and adversarial test MSEs drop, indicating that the overcompleteness of the learnable sparsifying transform promotes ADMM-DAD’s adversarial robustness. Of course, the MSEs increase as $\varepsilon$ also increases, but this is anticipated, since a stronger attack in the CS observations tantamounts to a DUN being less able to recover the data. Nevertheless, explaining the adversarial robustness of ADMM-DAD through the overcompleteness of the learnable sparsifier lays a fruitful ground, to identify “hyper-parameters” and properly fine-tune them, so as to boost adversarial robustness in the unfolding regime; we briefly mention this research line in Sec. 6.
@@
A note on the robustness gap. Based on Figures 1 and 3, we examine the robustness gap of ADMM-DAD, i.e., the difference between (19) and (20). We notice that the robustness gap slightly increases for the CIFAR10 with increasing $\varepsilon$ of the FGSM – albeit the gap’s scaling from one attack level to the next is reasonable – but the picture, e.g., Figure 1, is much better for the SVHN, where both test errors grow proportionally. This indicates that ADMM-DAD enjoys adversarial robustness to increasing attack levels. Especially in the case of $\varepsilon=0.01$ – which is still a non-negligible attack level – we deduce from Figure 3 that, for both datasets, the robustness gap is especially small, e.g., $\sim 10^{-4}$. Due to these interesting findings, explaining the robustness gap of DUNs, both from a theoretical and practical side, could be an inspiring future work (see also Sec. 6).
@@
Table 2: Comparison of ADMM-DAD – with overcompleteness $N=10n$ – to the baseline, both with 10 layers, against different FGSM attack levels $\varepsilon$, on CIFAR10 (top) and SVHN (bottom). Bold letters indicate the DUN that scores the best performance in terms of all metrics (19), (20), (21). Overall, ADMM-DAD outperforms the baseline, highlighting the advantage of overcompleteness over orthogonality in the unfolding regime.